OpenStudy (anonymous):

Assume that y varies inversely with x. If y = 7 when x = -3, find y when x = 7.

OpenStudy (minecrafter64):

17 I think

OpenStudy (anonymous):

\[\frac{ x _{1} }{ x _{2} }=\frac{y _{1} }{y_ 2 }\]

OpenStudy (anonymous):

that is the inverse variation formula..^^^ plug in the values you were given, so what is you're answer/

OpenStudy (anonymous):

\[\frac{ -3 }{ 7 }\frac{ 7 }{ y }\]

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

yeah, so what is y? And I hope I am right... someone correct me if I'm wrong!

OpenStudy (irishboy123):

Y = -21/x is but one of an infinite number of solutions to this. To solve the generalised solution y = a/x + b for a and b, you need 2 points.

OpenStudy (anonymous):

-21 ? the other answer choices i have are y = -7/3 and y= -3